OpenStudy (anonymous):

Where is the horizontal asymptote of the graph of y=3(2/3)^x-1

OpenStudy (anonymous):

@Nnesha

OpenStudy (jdoe0001):

for say any exponential function, with only "x" as exponent, where do they have their horizontal asymptote? like say \(\large 2^x\ or \ 3^x?\)

OpenStudy (anonymous):

The horizontalassymptote would be 2or 3

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

@jdoe0001 is helping you already :)

OpenStudy (jdoe0001):

you see the horizontal asymptote for that function, is the same as for any exponential function the only difference is the (x-1) <--- which is just a function transformation to be exact, is a horizontal shift from the "parent" function

OpenStudy (anonymous):

sois it 2/3

OpenStudy (jdoe0001):

hmm nope, that's just the base not the horizontal asymptote

OpenStudy (anonymous):

so it is -1

OpenStudy (jdoe0001):

notice all these exponential functions -> http://www.softmath.com/tutorials-3/relations/articles_imgs/2441/expone60.gif they all have the same horizontal asymptote though some are steeper than others the horizontal asymptote is the same anyhow.. need to dash but to get the transformation \(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards\\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ % template start \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ y = {\color{purple}{ A}} ( x + {\color{red}{ C}} ) + {\color{green}{ D}}\\ %template end \qquad\qquad \quad \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ {\color{red}{ C}}\end{array} \begin{array}{llll}{\color{red}{ C}} > 0 & to\ the\ left\\ {\color{red}{ C}} < 0& to\ the\ right\end{array} \end{array} }\)

OpenStudy (anonymous):

so the horizontal asymptote is 1